OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 7. Find the number.

OpenStudy (anonymous):

please show me the formula

OpenStudy (anonymous):

\[x+\frac{6}{x}=7\]

OpenStudy (anonymous):

6 and 1

OpenStudy (anonymous):

x + 6/x=7 x^2-7x+6=0 x=(-b+sqrt. b^2-4ac)/2a and the other value is (-b-sq.rt b^2-4ac)/2a

OpenStudy (anonymous):

this is sri dharacharya's method.hope this solve's ur problem.this is a direct method.

OpenStudy (anonymous):

\[x+\frac{6}{x}=7\] \[x^2+6=7x\] \[x^-7x+6=0\] \[(x-6)(x-1)=0\] \[x=6\] or \[x=1\]

OpenStudy (anonymous):

or by inspection